problem 1 Solve. 6x – 0.37 = 5(x – 0.28)
A) –0.69 B) –1.24 C) –0.89 D) –1.03

I thought it was d but i got it wrong

Answer by Nate(3500)

6x – 0.37 = 5(x – 0.28)
6x – 0.37 = 5x – 1.4
6x - 5x – 0.37 + 0.37 = 5x - 5x – 1.4 + 0.37
x = -1.03
Check:
6(-1.03) – 0.37 = 5(-1.03 – 0.28)
6(-1.03) – 0.37 = 5(-1.31)
-6.55 = -6.55
(D) is your solution.